The following information is available for Metlock Corporation for 2019 (its first year of operations).

1. Excess of tax depreciation over book depreciation, $40,000. This $40,000 difference will reverse equally over the years 2020–2023.
2. Deferral, for book purposes, of $19,800 of rent received in advance. The rent will be recognized in 2020.
3. Pretax financial income, $303,400.
4. Tax rate for all years, 20%.

Part 1

Compute taxable income for 2019.

Taxable income

$enter Taxable income in dollars

eTextbook and Media

List of Accounts

Part 2

Prepare the journal entry to record income tax expense, deferred income taxes, and income taxes payable for 2019. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Account Titles and Explanation

Debit

Credit

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

eTextbook and Media

List of Accounts

Part 3

Prepare the journal entry to record income tax expense, deferred income taxes, and income taxes payable for 2020, assuming taxable income of $297,100. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Account Titles and Explanation

Debit

Credit

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount
Answer #1
1
Pretax financial income 303400
Less: Excess of tax depreciation -40000
Add: Rent Received in advance 19800
Taxable income 283200
2
Account Titles and Explanation Debit Credit
Income tax expense 60680
Deferred tax asset 3960 =19800*20%
        Deferred tax liability 8000 =40000*20%
        Income taxes payable 56640 =283200*20%
3
Account Titles and Explanation Debit Credit
Income tax expense 61380
Deferred tax liability 2000 =8000/4
        Deferred tax asset 3960
        Income taxes payable 59420 =297100*20%